Mathematics 119, Fall Semester 1999-2000

Are you in the right course?

Math 119-120 is intended for students planning to enter pre-professional or biology programs. It is not intended for students who intend to major in engineering, physics, mathematics, or most of the chemistry programs. Those majors require Math 125-126.

Instructor: Professor Juan Migliore 364 CCMB (On Juniper, just south of the library) Phone: 631-7345

Office Hours: Tuesday 10:15–11:15 Wednesday 10:00-11:00 Or by appointment.

Tutorial Instructor: Michael Bergan 336 CCMB

TA’s Office Hours: Tuesday 3:00–4:00

Text: Single Variable Calculus (third edition), by James Stewart

Course Web Page: http://www.nd.edu/∼jmiglior/homepage119.html

Classes, Tutorials and Help Sessions

Class meets at 1:55 PM MWF in EART 102. Students are required to attend these classes.

Each student is also assigned to a Thursday tutorial section. It is your responsibility to know the time, place and section number of your tutorial section. At the tutorial, the previous week’s homework will be returned. Students will then be encouraged to ask questions about that homework or about the current homework. The purpose of the tutorial is to help students master the material currently being covered.

In addition to the tutorials, the professor and the tutorial instructor will have office hours each week at which you can get assistance in understanding the course work and doing the homework problems.

Examinations, homework and grades

There will be three one-hour examinations and one final examination (whose dates, times and locations are listed below). Each one-hour exam will be worth 100 points, and will be returned to you at the following tutorial session. The final exam is a two-hour exam and will be worth 150 points. The final exam will cover all the material of the course. There are a total of 450 possible points for the semester, and the cut-offs for the semester grades will be based on this total. Homework grades will be collected, and up to 5 “bonus” points will be added to your total from the exams.

Homework will be due at the Thursday tutorial and returned the following week. Usually three assignments will be due each Thursday; specific assignments due each week will be announced in that week’s lectures. Because it may happen that you have trouble with some homework problems and want another shot at them after you see the TA, we will accept homework as late as 4:00 of the following day (Friday). If you do not turn it in to the TA at the tutorial, you have to come to the Mathematics Department and put your homework in his mailbox (on the 3rd floor of CCMB).

The main purpose of collecting and returning homework is to let you know if you are doing the problems correctly. The homework grade (up to 5 “bonus” points) is designed to reward effort. Each problem is graded either 0 (if missing or complete nonsense) or 1 (for any honest attempt). So the total number of points on any assignment is simply the number of problems honestly attempted.

Both examinations and homework are conducted under the honor code. While cooperation in doing homework is permitted (and encouraged), copying is not. Exams are closed book and are to be done completely by yourself with no help from others. Calculators are not allowed on the exams.

A student who misses an examination will receive zero points for that exam unless he or she has written permission from the Dean of the First Year of Studies. (An excuse is almost certainly not going to be accepted if it is presented after the exam takes place.) Please be aware that travel plans are not considered to be a valid excuse by the Dean of the First year of Studies.

Exams

Date  Time  Place

Exam 1 Thursday, September 23, 1999 8:00 AM HAGR 117 Exam 2 Thursday, October 28, 19996 8:00 AM HAGR 117 Exam 3 Tuesday, November 30, 1999 8:00 AM STEP 100 Final Thursday, December 16, 1999 1:45 PM

(Location of final exam will be announced later.)

Mathematics 119, Fall Semester 1999-2000 HOMEWORK ASSIGNMENTS

Number Assignment
1  p. A15 # 7, 8, 21–23, 25–28, 31–33, 37–40
2  p. 14 # 2–4, 11–14, 20–22, 27, 31, 32, 41, 43, 44, 51–53
3  p. A22 # 1–8
p. A15 # 1–4
4  p. 50 # 2–5
5  p. 59 # 3–5, 7–9
6  p. 68 # 1, 15–20, 23, 24, 26, 27, 29, 31, 32
7  p. 87 # 1, 2, 32–34, 37–39
8  p. 96 # 1, 4, 5, 6, 11, 12
9  p. 109 # 1, 3–5, 8, 13–15, 23,24,26
10  p. 120 # 1, 2, 4, 10–13, 23, 24, 34, 36–39, 59, 60
11  p. A30 # 1, 3, 7, 9, 13–15, 23, 27, 28, 30, 32 ,42–46, 53
12  p. A30 # 65, 66, 69
p. 137 # 1, 3, 5–8, 10–14, 16
13  p. 137 # 20–22, 24, 25, 30, 32, 34, 37
14  p. 144 # 10, 11, 13, 17, 21–26, 35, 37, 38 49, 50, 57, 59, 60
15  p. 150 # 5, 8, 9, 13, 14, 16, 25
16  p. 155 # 3–6, 14, 18, 20, 27, 43, 44
17  p. 130 # 7, 8, 13, 15, 23, 24
p. 155 # 36, 37, 40, 41
18  p. 160 # 1, 3, 7–12
19  p. 160 # 16–18, 25, 26, 29
20  p. 188 # 21–24, 31–33, 37, 38, 41–43, 47–51
21  p. 199 # 3, 4, 12, 16, 18, 20, 22, 23, 25, 26
22  p. 205 # 1, 4, 7, 8, 13, 14, 15, 17, 21– 24
23  p. 216 # 9–14, 24, 36, 37
24  p. 224 # 4, 5, 15, 16, 39
25  p. 236 # 1, 3, 6–12
26  p. 236 # 13, 14, 22, 27, 28, 29
27  p. 249 # 2, 5, 8–10, 13, 15, 16, 27, 28 , 35, 50, 53, 55
28  p. 262 # 2, 12, 13, 19, 21, 26, 29, 30
29  p. 271 # 1, 2, 3, 23, 24
30  p. 280 # 7, 9, 15, 16, 23, 24, 26
31  p. 291 # 5, 6, 9, 10–13, 17, 18, 23, 32, 46, 51, 52
32  p. 291 # 63–66, 69–72, 74, 79
33  p. 300 # 1, 3, 5, 11, 13